OpenStudy (anonymous):

Another question about projectiles launched at an angle. A baseball is thrown at an angle of 25 degrees relative to the ground at a speed of 23.0m/s. If the ball was caught 42.0m from the thrower, how long was it in the air? How high above the thrower did the ball travel?

OpenStudy (anonymous):

\[t=\frac{ 42 }{ 23\cos25 }\]

OpenStudy (anonymous):

y is now the height of the tower\[y=t(Vsintheta -4.9t)\]

OpenStudy (anonymous):

where did you get the -4.9?

OpenStudy (anonymous):

it's 1/2 * a since a is -g and g is 9.8 acceleration in the y direction is -g for projectile then 1/2 of -9.8 = -4.9

OpenStudy (anonymous):

Okay thanks. duh haha.

OpenStudy (anonymous):

good lol

OpenStudy (anonymous):

so basically use Vy=vi*sin25-1/2gt ???

OpenStudy (anonymous):

yea but remember it's vi*sin25*t-1/2*g*t^2

OpenStudy (anonymous):

i got a physics test tomorrow so i'm practicing by answering question

OpenStudy (anonymous):

same here I'm totally flipping out

OpenStudy (anonymous):

im doing college phsyics , harder stuff lol

OpenStudy (anonymous):

ugh I didn't get the right answer...it's supposed to be 4.8 m I got the seconds just not the Dy yet

OpenStudy (anonymous):

and I got -8.4m o_o

OpenStudy (anonymous):

dy is 42 we assume the projectile is launched from the origin of our coordinate system

OpenStudy (anonymous):

so wait. what am I solving for when it asks how high above the thrower did the ball travel?

OpenStudy (anonymous):

vi is v initial so h=(vi^2*sin^2thetha)/2g that's to find height, which is gonna be the max height of the parabola

OpenStudy (anonymous):

 just to be clear, is this drawing sort of right? for a normal diagram I guess

OpenStudy (anonymous):



OpenStudy (anonymous):

still not getting the height :/